### v1

# Diff3DS: Generating View-Consistent 3D Sketch via Differentiable Curve Rendering

Abstract
3D sketches are widely used for visually representing the 3D shape and structure of objects or scenes. However, the creation of 3D sketch often requires users to possess professional artistic skills. Existing research efforts primarily focus on enhancing the ability of interactive sketch generation in 3D virtual systems. In this work, we propose Diff3DS, a novel differentiable rendering framework for generating view-consistent 3D sketch by optimizing 3D parametric curves under various supervisions. Specifically, we perform perspective projection to render the 3D rational Bézier curves into 2D curves, which are subsequently converted to a 2D raster image via our customized differentiable rasterizer. Our framework bridges the domains of 3D sketch and raster image, achieving end-to-end optimization of 3D sketch through gradients computed in the 2D image domain. Our Diff3DS can enable a series of novel 3D sketch generation tasks, including text-to-3D sketch and image-to-3D sketch, supported by the popular distillation-based supervision, such as Score Distillation Sampling (SDS). Extensive experiments have yielded promising results and demonstrated the potential of our framework. Project: https://yiboz2001.github.io/Diff3DS/
Figure 1: In this paper, we propose Diff3DS, a novel differentiable rendering framework for generating view-consistent 3D sketch from flexible inputs such as a single image or text.
1 Introduction
3D sketches, which utilize strokes to emphasize abstraction and visually encapsulate the 3D shape and structure of objects or scenes, serve as indispensable tools for visualizing concepts and ideas. Existing research efforts have primarily focused on exploring how to enable interactive creation of 3D sketch in 3D virtual systems [44, 12, 5] and immersive environments [53, 48, 19, 17, 18, 54, 4, 55]. Such interactive 3D sketch creation tools provide artists with unprecedented freedom, allowing them to draw their inspirations directly in the immersive 3D environment. However, it usually requires users to possess professional artistic skills and experience for 3D operation, making the 3D sketch creation less friendly for ordinary users. There is a notable absence of a user-friendly 3D sketch generation method in the community.
In recent years, content reconstruction or generation via differentiable rendering has gained great attention in the fields of computer graphics and vision research. The goal is to convert the traditional rendering pipeline into a differentiable image synthesis process and employ the 2D image supervision to optimize the object or scene representation. For example, NeRF [24] bridges the 3D neural representation and raster image domain with a differentiable volume rendering pipeline, and has contributed to the further development of the 3D generation [25, 38, 15, 28, 21] by leveraging the continuous advancements in multimodal supervision [32, 28]. For the generation of vectorized or parametric content, differentiable rendering has also been investigated [20, 33, 10, 16, 49, 31, 50, 56, 6, 40, 39]. One representative work is DiffVG [20], which proposes a differentiable 2D vector graphics rasterizer to compute gradients from raster images and optimize the vectorized image (e.g., 2D Bézier curves and other parametric shapes) via back-propagation. For generation of sketch, CLIPasso generates 2D sketches using a differentiable rasterizer to optimize the parameters of the Bézier curves directly with respect to a CLIP-based perceptual loss. Recently, 3Doodle [8] generates view-consistent 3D sketches of the target object by directly optimizing the parameters of 3D strokes to minimize the perceptual losses for given multi-view images.
Although 3Doodle achieves promising results for generating 3D sketches, it has two main limitations. First, 3Doodle performs orthogonal projection to obtain the 2D sketch rendering and ignores the depth order of the curves, which may lead to errors in the 3D sketch. Second, its optimization requires supervision from multiview images, which makes it more like a reconstruction pipeline instead of a more flexible generation pipeline. In this paper, we aim to solving above issues and present Diff3DS, a novel differentiable rendering framework for generating view-consistent 3D sketch from flexible inputs such as a single image or text. Specifically, we represent the 3D sketch as a set of 3D rational Bézier curves and perform the perspective projection to obtain the 2D rational Bézier curves. Then, we propose a new differentiable rasterizer based on DiffVG [20] to accurately render the projected 2D curves based on the depth order. Our framework supports end-to-end optimization of 3D curve primitives under flexible supervisions such as the distillation-based loss. By employing the recent Score Distillation Sampling (SDS) [28, 22] to distill prior knowledge from pre-trained 2D image generation model, we achieve 3D sketch generation from the text or single image input. We conduct quantitative and qualitative comparisons between Diff3DS and related methods, and the results show the superiority of our approach in generating view-consistent 3D sketch. Ablation studies also verify the effectiveness our our key hyperparameters and components.
We summarize our contributions as follows: (1) We propose Diff3DS, a novel differentiable rendering framework for generating view-consistent 3D sketch from flexible inputs such as a single image or text. (2) We are the first to represent the 3D curve as 3D rational Bézier curve and design a new differentiable rasterizer to render the perspectively projected 2D rational Bézier curves, while considering the depth order. (3) We conduct comprehensive experiments on novel text-to-3D and image-to-3D sketch generation tasks and the results demonstrate the superiority of our method.
2 Related Works
3D Sketch Generation
Existing research on 3D sketch generation mainly focus on interactive artistic sketch creation within 3D virtual systems [44, 12, 5], and immersive environments [53, 48, 19, 17, 18, 54, 4, 55]. However, these works usually require users to possess professional artistic skills and 3D operation experience, making it less user-friendly for ordinary users. The recent 3Doodle [8] is able to generate expressive 3D sketches from multiview image observations, but it is more like multiview reconstruction instead of generation and the quality of the generated results highly depends on the number of observation viewpoints. By distilling prior knowledge from pre-trained 2D image generation models using methods like Score Distillation Sampling [28, 22], we achieve user-friendly and view-consistent 3D sketch generation from flexible text or single image input.
Differentiable Rendering of Curves
There have been several works for differentiable rendering of curves. DiffVG [20] proposes a differentiable rasterizer for the creation of vector graphics represented by 2D curves and shapes. DRPG [46] investigates differentiable rendering of 3D parametric curves and surfaces, by piecewisely approximating the continuous parametric representations with a triangle mesh. The 3Doodle [8] shares a similar differentiable rendering pipeline with us, including projecting the 3D Bézier curves to 2D Bézier curves, which are then rendered by the differentiable rasterizer from DiffVG. However, 3Doodle performs orthogonal projection to obtain 2D Bézier curves and ignores the depth ordering between the curves. Instead, we perform perspective projection to the 3D rational Bézier curves and designs a new DiffVG-based ratersizer which can maintain the depth ordering between the projected 2D rational Bézier curves.
Multimodal-Driven 3D Content Generation
Motivated by the success of the text-to-image research, pioneering works [28, 41] introduce the Score Distillation Sampling (SDS) algorithm and leverage the pre-trained text-to-image models as prior knowledge to optimize 3D representations. Subsequent works combine SDS with various differentiable 3D representations to explore the capabilities of text-to-3D object generation, such as DMTet [21, 7] and 3D Gaussians [37, 52]. Furthermore, Zero-1-to-3 [22] proposes an image-guided SDS algorithm that distills the 3D-consistent priors to optimize 3D representations given the input image, and its pipeline inspires the development of subsequent image-to-3D object methods [22, 30, 42, 47]. Distinguishing from prior endeavors, our work stands out by focusing on generating view-consistent 3D sketch by optimizing 3D parametric curve primitives.
3 Preliminary
DiffVG [20] proposes a differentiable rasterizer for vector graphics that supports converting 2D Bézier curves to the raster image domain and back-propagating the gradient for optimization.
Given the 2D curves parameters $\Theta$, the vector graphics scene is defined as $f(x,y,\Theta)$, and the raster image is defined as the 2D grid sampling over the space of $f(x,y,\Theta)$. To compute the color of a 2D location $(x,y)\in\mathbb{R}^{2}$, DiffVG utilizes the inside-outside test [26] to find the curves overlapped with the location initially. Subsequently, it sorts them according to a user-specified order and calculates the color using alpha blending [29]. Due to the inside-outside test, the scene function $f$ is not differentiable with respect to curve parameters. Thus, it uses the anti-aliasing technique to make the pixel color differentiable. By prefiltering $f$ over a convolution kernel $k$ with support $A$, sampling $I(x,y)$ at pixels can yield an alias-free image:
[…]
Due to the integration over the filter support region, the focus shifts from the color value at the center point to the average pixel color. The continuous change in the average color induced by the curve movements makes the function $I(x,y)$ differentiable.
The goal is to compute the gradients of $I$ with respect to $\Theta$:
[…]
DiffVG proposes two approaches Monte Carlo Sampling and Analytical Prefiltering to evaluate the pixel integral, which does not have a closed-form solution in general. We focus on the first approach which performs better. The pixel integral can be discretized with Monte Carlo sampling [11]:
[…]
but geometric discontinuities hinder the interchangeability of the integral and differential operators, which consequently obstructs the direct representation of $\frac{\partial I}{\partial\Theta}$ as a discretizable integral. Thus, it rewrites the pixel color as the sum of integrals over multiple sub-regions $A_{i}$, and applies the Reynolds transport theorem to handle all the discontinuous changes on the boundary:
[…]
where $g(u,v)$ is the multiplication of the scene function $f$ and kernel $k$ for brevity, $\partial A_{i}$ is the boundary of area $A_{i}$, $n(t)$ is the outward normal of $\partial A_{i}$, and $p(t)$ denotes the 2D points on the $\partial A_{i}$. The first integral is responsible for the differentiation of color and the transparency, while the second term is responsible for the change of the boundaries, and the boundary integral can be estimated by the Monte Carlo estimator as:
[…]
where $\epsilon$ is a small number and $P(p_{j}|c)P(c)$ denotes the probability density of sampling the curve $c$ and point $p_{j}$ on the boundaries.
4 Diff3DS
4.1 Overview
We propose Diff3DS, a novel differentiable rendering framework for generating view-consistent 3D sketch by optimizing 3D parametric curves. A 3D sketch $\tilde{\Theta}$ is defined as a collection of 3D parametric curve strokes that reside in 3D space. We particularly focus on the 3D rational Bézier curve and our derivation supports the 3D linear, quadratic, and cubic rational Bézier curves. The theoretical trainable parameters include the control point position, curve width, and rgb-alpha color. Follow the previous methods [40, 39, 8], we represent each stroke by a 3D cubic Bézier curve and only optimize the control point position in our current implementation. Moreover, weights for rational Bézier curves are disregarded to simplify the computation. Fig. 2 illustrates the pipeline.
Diff3DS is designed with three stages. First, to render the 3D rational Bézier curve strokes $\tilde{\Theta}$ at a given viewpoint, we perspectively project these curves onto the camera plane (Sec. 4.2). Then, we render the projected 2D rational Bézier curves $\Theta$ to the raster image using a customized differentiable rasterizer. (Sec. 4.3). Finally, we will perform back-propagation with the gradient computed from the rendered image to optimize the 3D sketch parameters (Sec. 4.4).
4.2 3D Rational Bézier Curve Projection
In this section, we mainly explain how to perspectively project the 3D rational Bézier curves in camera coordinates onto the 2D camera focal plane. Given the control points $\tilde{P}_{i}=(\tilde{P}_{i}^{x},\tilde{P}_{i}^{y},\tilde{P}_{i}^{z})\in\mathbb{R}^{3}$, the associated non-negative weights $\tilde{w}_{i}$ and the Bernstein polynomials $B_{i,n}(t)$, for $0\leq t\leq 1$, define the 3D rational Bézier curve $\tilde{p}(t)$ of degree $n$ by:
$\tilde{p}(t)$ $=\sum_{i=0}^{n}\frac{B_{i,n}(t)\tilde{w}_{i}}{\sum_{j=0}^{n}B_{j,n}(t)\tilde{w}_{j}}\tilde{P}_{i}=(\tilde{x}(t),\tilde{y}(t),\tilde{z}(t)).$ (6)
Assume the focal plane is $z=f$, where $f$ is the focal length, the corresponding projected 2D curve $p(t)$ under the pinhole camera perspective projection can be defined as:
$p(t)=(x(t),y(t))=(\frac{\tilde{x}(t)}{\tilde{z}(t)}f,\frac{\tilde{y}(t)}{\tilde{z}(t)}f).$ (7)
Due to rational Bézier curves are projective invariant, the curve $p(t)$ still is a 2D rational Bézier curve defined by projected control points. Given projected 2D control points $P_{i}=(P_{i}^{x},P_{i}^{y})=(\frac{\tilde{P}_{i}^{x}}{\tilde{P}_{i}^{z}}f,\frac{\tilde{P}_{i}^{y}}{\tilde{P}_{i}^{z}}f)\in\mathbb{R}^{2}$ and adjusted weights $w_{i}=\tilde{w}_{i}\tilde{P}_{i}^{z}$, $p(t)$ can be written as following:
$p(t)=\sum_{i=0}^{n}\frac{B_{i,n}(t)w_{i}}{\sum_{j=0}^{n}B_{j,n}(t)w_{j}}P_{i}=(x(t),y(t)).$ (8)
We provide the detailed proof in Appendix B.
Figure 2: We generate the 3D sketch $\tilde{\Theta}$, represented as a set of 3D strokes, from the text or image input. We render the raster image $I$ from the random camera via our differentiable renderer (Sec. 4). Then, the pre-trained diffusion model, conditioned on the input, diffuses the rendering $I$ and predicts the pseudo ground truth $\hat{I}_{0}$. The discrepancies between $\hat{I}_{0}$ and $I$ are used to update the 3D sketch.
4.3 2D Rational Bézier Curves Differentiable Rasterization
This section primarily focuses on the differentiable rasterization of 2D rational Bézier curves $\Theta$ obtained through projection.
To render the projected curves and maintain their depth orders, we customize a new differentiable rasterizer based on DiffVG [20]. Given a 2D pixel location $(x,y)$ and projected curves $\Theta$, we employ the inside-outside test to identify all the overlapping curves with the location. The inside-outside test identifies a curve as overlapping with a location if the closest distance between them is less than half of the curve width, and the closest point on each overlapping curve with the location are considered as the overlapping point. Then, we sort the overlapping point to calculate the color of the pixel location $(x,y)$ using alpha blending [29]. Our sort order follows the depth order between curve primitives. We employ the unprojected z-depth in the camera space of the overlapping point as the primary sort order. To mitigate the z-fighting issue [^1] https://en.wikipedia.org/wiki/Z-fighting caused by floating-point error, we also employ a user-specified order as the secondary order.
Identifying the overlapping curve is challenging. For the $n$th-degree rational curve $p(t)$, computing the $t$ that minimizes the distance $(p(t)-q)^{2}$ where $q=(x,y)$ equals to solve the roots of a polynomial with degree $23n-$ 2. For the cubic rational curve, a 7th order polynomial needs to be solved, which doesn’t have a closed-form solution. Inspired by DiffVG, we solve the polynomial using bisection and the Newton-Raphson method [45]. The iterative solver obtains its initial guess from isolator polynomials [34] – the real roots of a 7th order polynomial can be isolated by the roots of two cubic auxiliary polynomials. Please refer to Appendix C for more details.
After obtaining the 2D curve scene function $f(x,y,\Theta)$, the rendered raster image $I$ will be calculated using the Monte Carlo Sampling strategy (Sec. 3) with Eq. 3. Fig. 3 shows an example.
Figure 3: A rasterized result contains 3 quadratic rational Bézier curves and a line. All curves share the same control point positions in pixel space and the depths but with different weights. Our rasterizer accurately renders the curves and faithfully maintains occlusions according to the depth order. The user-specified sort order results in the pixel at position (1023, 0) being the color of Curve C, as it has a higher priority than the other overlapping curve points with the same depth.
4.4 Gradients Back-propagation
The rendering framework is fully differentiable since the projection and rasterizer are differentiable. Thus, the gradients of $I$ with respect to 3D curves parameter $\tilde{\Theta}$ can be calculated with the chain rule: $\frac{\partial I(x,y)}{\partial\tilde{\Theta}}=\frac{\partial I(x,y)}{\partial\Theta}\frac{\partial\Theta}{\partial\tilde{\Theta}},$ where $\frac{\partial I}{\partial\Theta}$ denotes the gradients of $I$ with respect to 2D projected curves $\Theta$ that can be computed using Eq. 4 and Eq. 5, and $\frac{\partial\Theta}{\partial\tilde{\Theta}}$ denotes the gradients of projected 2D curves $\Theta$ with respect to 3D curves $\tilde{\Theta}$. $\frac{\partial I(x,y)}{\partial\Theta}$ calculates the gradients of control point positions, color, and width for projected 2D curves. The position gradients of 3D curve control points are calculated with $\frac{\partial\Theta}{\partial\tilde{\Theta}}$, which is based on the gradients of the projected 2D control points. The color and width gradients of 3D curves directly inherit the corresponding gradients of the projected 2D curves.
5 Multimodal-Driven 3D Sketch Generation
Diff3DS supports end-to-end optimization of 3D sketch under flexible supervisions. By employing the Score Distillation Sampling algorithm, we propose the promising tasks of text-to-3D sketch and image-to-3D sketch, with the goal of generating 3D sketch from flexible text or single image input.
5.1 Text-to-3D Sketch
Pioneering works [28, 41] propose the Score Distillation Sampling (SDS) algorithm, utilizing the prior of a pre-trained 2D text-to-image model to optimize 3D representations. By integrating SDS with Diff3DS, we distill the diffusion prior to generate the 3D sketch $\tilde{\Theta}$ from the input text. Given a pre-trained diffusion model $\epsilon_{\phi}$, text embedding $y$, rendered image $I$ with our rasterizer $\mathcal{R}$ and the noise timestep $t$, image $I$ will be added with noise to obtain a noisy image $I_{t}$: $I_{t}=\sqrt{\bar{\alpha}_{t}}I+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$ where $\epsilon\sim\mathcal{N}(0,I)$ and $\bar{\alpha}_{t}$ is the cumulative product of scaling at timestep $t$. Then the gradient of SDS loss on $\tilde{\Theta}$ is given by:
$\nabla_{\tilde{\Theta}}\mathcal{L}_{SDS}(\phi,I=\mathcal{R}(\tilde{\Theta}))=\mathbb{E}_{t,\epsilon}\left[\omega(t)(\hat{\epsilon}_{\phi}(I_{t};y,t)-\epsilon)\frac{\partial I}{\partial\tilde{\Theta}}\right],$ (9)
where $\omega(t)$ is a weight based on the timestep $t$, and the predicted noise sampled by classifier-free guidance (CFG) [13] with weight $\lambda$ denotes as:
$\hat{\epsilon}_{\phi}(I_{t};y,t)=\epsilon_{\phi}(I_{t};\emptyset,t)+\lambda(\epsilon_{\phi}(I_{t};y,t)-\epsilon_{\phi}(I_{t};\emptyset,t)).$ (10)
By deriving the pseudo ground truth image $\hat{I}_{0}$ with one-step denoising:
$\hat{I}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(I_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{\epsilon}_{\phi}(I_{t};y,t)),$ (11)
we can transform Eq. 9 into an equivalent form as follows:
$\hskip 1.42262pt\nabla_{\tilde{\Theta}}\mathcal{L}_{SDS}(\phi,I=\mathcal{R}(\tilde{\Theta}))$ $=\mathbb{E}_{t,\epsilon}\left[\omega(t)\frac{\sqrt{\bar{\alpha}_{t}}}{1-\sqrt{\bar{\alpha}_{t}}}(I-\hat{I}_{0})\frac{\partial I}{\partial\tilde{\Theta}}\right],$ (12)
which can be regarded as matching the input view $I$ with predicted $\hat{I}_{0}$.
Similar with [28], due to the bias of 2D text-to-image model, we encounter serious view inconsistency problem known as the Janus problem. This problem persists even when directional prompts (e.g., "the front view of…") are used to distinguish between different views. To mitigate this issue, we integrate MVDream [35] with our framework, which is able to generate 3D-aware four view images using input text and camera poses.
Figure 4: Pseudo ground truth $\hat{I}_{0}$ visualization. As the CFG weight $\lambda$ decreases, the effective supervision region of $\hat{I}_{0}$ also decreases.
The impact of the CFG weight $\lambda$ on the optimization process has received attention in existing text-to-3D methods. By visualizing the pseudo ground truth $\hat{I}_{0}$ after denoising with different levels of weight $\lambda$ at large noise timesteps in Fig. 4, we find that as $\lambda$ decreases, the effective supervision region of $\hat{I}_{0}$ decreases, hereby impacting the final shape of results.
5.2 Image-to-3D Sketch
Zero-1-to-3 [22] proposes an image-guided SDS method that conditions on the input view image $\tilde{I}$ and relative camera extrinsic ($R$, $T$), based on its view-conditioned model. Our framework is able to generate 3D sketch $\tilde{\Theta}$ from the input reference image $\tilde{I}$ by distilling the 3D consistent prior:
$\nabla_{\tilde{\Theta}}\mathcal{L}_{SDS}(\phi,I=\mathcal{R}(\tilde{\Theta}))=\mathbb{E}_{t,\epsilon}\left[\omega(t)(\hat{\epsilon}_{\phi}(I_{t};\tilde{I},R,T,t)-\epsilon)\frac{\partial I}{\partial\tilde{\Theta}}\right].$ (13)
5.3 Dynamic Noise Deletion
We find that our framework is affected by the issue of gradient sparsity in the Monte Carlo Sampling strategy. In Reynolds’ formula, the gradient from pixel space is determined by the curve boundaries. This results in the loss of information from pixels that are not incident to those boundaries. In cases where significant displacements of curves are required, the optimization tends to get trapped in a local optimum state: curves contract the boundaries to minimize the loss function. During our training process, some curves tend to contract to extremely small lengths as noises. To obtain clean results, we propose the Dynamic Noise Deletion strategy. We discretize each curve into $n$ line segments ($n=20$) and accumulate their lengths as the total length of the curve. The curve with lengths below a specified threshold will be considered as noise and removed dynamically.
5.4 Time Annealing Schedule
To further enhance the performance of the results, we employ a time annealing schedule similar with [43, 57, 35, 14]. During the optimization process, we gradually adjust the maximum and minimum time step for SDS in a linear manner. The large noise timestep focuses on aligning the curves with the semantic content of the input text or image during the initial training phase. Conversely, the small noise timestep employed during the later training phase focuses on further enhancing the details.
Figure 5: Qualitative results of the text-to-3D sketch task. Existing text-to-3D methods fail to generate sketch-style results, even with the addition of the "sketch in black and white, line drawing" suffix.
6 Experiments
6.1 Implementation Details
We implement our rendering framework in C++/CUDA with a PyTorch interface [27], and follow the STEP standard (ISO 10303-21) [23] to format and save our 3D curve results. In the experiment, a user-specified number of curves will be randomly initialized within a sphere of radius 1.5. We randomly sample the camera position using the radius from 1.8 to 2.0, with the azimuth in the range of -180 to 180 degrees, the elevation in the range of 0 to 30 degrees and the field of view (fov) of 60 degrees. Notice that Stable-Zero123 abnegates the relative distance for simplifying, we fix its radius at $2.0$. For the pre-trained model, we apply Stable Diffusion 2.1 [2] and MVDream [35] for the text-to-3D sketch task, with the CFG weight of $50$. And we apply Zero-1-to-3 [22] and Stable-Zero123 [3] for the image-to-3D sketch task, with the CFG weight of $7.5$ followed [22, 3] . For all tasks, the total number of training steps is 4000. Starting from step 2000, we dynamically delete the noise every 100 steps. The training process requires 1 hour for the text-to-3D sketch task and 2 hours for the image-to-3D sketch task on a single NVIDIA A10 GPU with a batch size of 4.
6.2 Text-to-3D Sketch
Baselines
To the best of our knowledge, we are the first text-to-3D sketch method. We select the existing text-to-3D object methods DreamGaussian [37] and MVDream [35] as compared baselines, and further choose the text-to-2D sketch method DiffSketcher [49] as the additional reference.
Quantitative Comparisons
We collect 35 text prompts from previous works [28, 35] and websites. To encourage baseline methods to generate sketch-style results, we append the suffix "sketch in black and white, line drawing" to each prompt inputted to the baselines. We measure the CLIP [32] text-image similarity (CLIP-ScoreT) metric to evaluate the semantic coherence between the input text prompt and the rendered images. For each generated 3D content, we compute the average metric across four orthogonal view images relative to the input text prompt. Table 2 shows the quantitative results of the text-to-3D sketch task. DiffSketcher achieves the best performance, while our method outperforms all the text-to-3D baselines.
Qualitative Comparisons
Fig. 5 shows the qualitative results of the text-to-3D sketch task. We compared the 3D results of the text-to-3D baseline method with our method, and further provide the 2D results of DiffSketcher. Compared text-to-3D baselines utilize Stable Diffusion [2] as the prior provider but do not fully inherit its capability. They struggle to generate the sketch-style content, even with the additional suffix. Our method and DiffSketcher can generate the 3D sketch and 2D sketch similar with the input text.
Table 1: Quantitative results of the text-to-3D sketch task. The CLIP-ScoreT metric is shown.
Method CLIP-ScoreT $\uparrow$
ViT B/16 ViT B/32
DiffSketcher [49] 33.64 33.98
DreamGaussian [37] 26.25 25.12
MVDream [35] 28.39 27.08
Our 30.35 30.00
Table 2: Quantitative results of the image-to-3D sketch task. The CLIP-ScoreI metric is shown.
Method CLIP-ScoreI $\uparrow$
ViT B/16 ViT B/32
CLIPasso [40] 83.79 82.72
NEF [51] 66.69 66.53
Our 67.81 68.54
6.3 Image-to-3D Sketch
Baselines
To the best of our knowledge, we are the first to generate 3D sketch from a single image. The closest works to ours are Neural Edge Fields (NEF) [51] and 3Doodle [8], which generate 3D curves from multiview images. However, since 3Doodle has not released the code, we select NEF for comparison. NEF reconstructs 3D curves from the 2D multiview edge maps estimated from PidiNet [36]. We also choose the image-to-2D sketch method CLIPasso [40] as the additional reference.
Quantitative Comparisons
We collect 25 images generated with Imagine [1], and measure the CLIP-Score [32] visual similarity (CLIP-ScoreI) metric to evaluate the semantic coherence between the input reference image and the rendered images. For each generated 3D content, we render four orthogonal view images to calculate the average metrics with the reference image. Table 2 illustrates that CLIPasso performs the best, while our method outperforms NEF.
Figure 6: Qualitative results of the image-to-3D sketch task. NEF fails to exact reasonable 3D curves from the input image.
Qualitative Comparisons
Fig. 6 shows the qualitative results of the image-to-3D sketch task. As a reconstruction method, NEF fails to extract reasonable 3D curves from a single input image because it lacks the ability to predict the unobserved regions. Our method and CLIPasso can generate the 3D sketch and 2D sketch, respectively, conditioned on the input image.
6.4 Ablations and Analysis
Figure 7: Illustration of the optimization process.
Optimization Process Visualization
Fig. 7 illustrates the visualization of the training optimization process. The optimization starts with randomly initialized curves. As time progresses, curves gradually optimize to resemble the input text or image. Dynamic Noise Deletion intervenes during the middle stage of training to remove the noise generated during the training process.
Effect of Hyperparameters
Fig. 8 illustrates the effect of hyperparameters. Fig. 8 (a) illustrates the effect of different CFG weights $\lambda$ in the text-to-3D sketch task. A larger weight leads to a stronger shape prior, while a smaller weight results in shape degradation. Fig. 8 (b) illustrates the effect of various pre-trained models. The Janus Problem easily arises when using Stable Diffusion, and MVDream partially mitigates this issue (e.g., the bicycle with three wheels in the bird’s-eye view). Then Zero-1-to-3 is not sufficiently 3D consistent in certain examples (e.g., the upper part of the scissors is bent in the side view). Fig. 8 (c) illustrates that increasing the initial curve number can enhance the details of results, even though the final curve number is automatically determined by Dynamic Noise Deletion. Too few curves may fail to represent a complete 3D object.
(a) CFG weight.
(b) Pre-trained Models.
(c) Initial Curve Number.
Figure 8: Effect of hyperparameters.
Ablation Study of Designed Components
We evaluate the contributions of the Time Annealing Schedule and Dynamic Noise Deletion components, as shown in Fig. 9. The omission of Time Annealing Schedule affects the conformity of the generated results with the semantic meaning of the input text or the level of detail in the reference image, and the omission of Dynamic Noise Deletion leads to the retention of some noise in the results.
Figure 9: Ablation study on Time Annealing Schedule (TAS) and Dynamic Noise Deletion (DND).
7 Conclusions
In this paper, we propose Diff3DS, a novel differentiable rendering framework for generating view-consistent 3D sketch by optimizing 3D parametric curves under various supervisions. By employing the recent Score Distillation Sampling (SDS) to distill prior knowledge from pre-trained 2D image generation model, we achieve 3D sketch generation from the flexible text or single image input. Extensive experiments have demonstrated the potential of our framework.
Limitations and Future Work
First, Diff3DS inherits the sparse gradient issue from DiffVG and can only optimize continuous parameters. Moreover, our downstream task assumption does not distinguish between the view-independent curve (e.g., feature lines) and the view-dependent curve (e.g., contours of smooth surface boundaries), limiting the expressive capacity of the overall shape of a 3D object, which is extensively discussed in 3Doodle. Incorporating the diverse curve representations of 3Doodle is a promising solution. Also, one future direction is to extend our method for scene-level generation task, such as text-to-3D scene sketch.
Supplementary
Overview
This supplementary material is structured into multiple sections that offer further details and analysis pertaining to our Diff3DS research. Concretely, it will hone in on the topics that follow:
  • $\bullet$
    In section A,we provide the implementation details of Diff3DS.
  • $\bullet$
    In section B, we provide a detailed proof for the 3D rational Bézier curve projection.
  • $\bullet$
    In section C, We provide a detailed analysis for finding the closest distance between a point and a rational Bézier curve.
  • $\bullet$
    In section D, we provide more high quality sketch results.
Appendix A Implementation Details
In the experiment, a user-specified number of curves will be randomly initialized within a sphere of radius 1.5. The default curve number is set to 56. We randomly sample the camera position using the radius from 1.8 to 2.0, with the azimuth in the range of -180 to 180 degrees, the elevation in the range of 0 to 30 degrees and the field of view (fov) of 60 degrees. Notice that Stable-Zero123 abnegates the relative distance for simplifying  [^2] https://github.com/threestudio-project/threestudio/issues/360, we fix its radius at $2.0$.
For the pre-trained model, we apply Stable Diffusion 2.1 [2] and MVDream [35] for the text-to-3D sketch task, with the CFG weight of $50$. And we apply Zero-1-to-3 [22] and Stable-Zero123 [3] for the image-to-3D sketch task, with the CFG weight of $7.5$ followed by [22, 3] .
For all tasks, the total number of training steps is 4000. Starting from step 2000, we dynamically delete the noise every 100 steps. The training process requires 1 hour for the text-to-3D sketch task and 2 hours for the image-to-3D sketch task on a single NVIDIA A10 GPU with a batch size of 4. To optimize the control point positions, we use the Adam optimizer and set the learning rate of the optimizer to 0.002. For the time annealing schedule, we prefer to decrease the maximum and minimum time steps from 0.85 to 0.3 and 0.1, respectively, over the first 3600 steps. Due to limited resources, not all possible combinations of the hyper-parameters related to the time steps have been fully explored, and there may be other configurations that could produce better results.
Appendix B 3D Rational Bézier Curve Projection
Given the control points $\tilde{P}_{i}=(\tilde{P}_{i}^{x},\tilde{P}_{i}^{y},\tilde{P}_{i}^{z})\in\mathbb{R}^{3}$, the associated non-negative weights $\tilde{w}_{i}$ and the Bernstein polynomials $B_{i,n}(t)$, for $0\leq t\leq 1$, define a $n$th-degree 3D rational Bézier curve $\tilde{p}(t)$ by:
$\tilde{p}(t)$ $=\sum_{i=0}^{n}\frac{B_{i,n}(t)\tilde{w}_{i}}{\sum_{j=0}^{n}B_{j,n}(t)\tilde{w}_{j}}\tilde{P}_{i}=(\tilde{x}(t),\tilde{y}(t),\tilde{z}(t)).$
Assume the focal plane is $z=f$, where $f$ is the focal length, the corresponding projected 2D curve $p(t)$ under the pinhole camera perspective projection can be defined as:
$p(t)=(x(t),y(t))=(\frac{\tilde{x}(t)}{\tilde{z}(t)}f,\frac{\tilde{y}(t)}{\tilde{z}(t)}f).$
Due to rational Bézier curves are projective invariant, the 2D curve $p(t)$ still is a rational Bézier curve defined by projected control points. Given projected 2D control points
$P_{i}=(P_{i}^{x},P_{i}^{y})=(\frac{\tilde{P}_{i}^{x}}{\tilde{P}_{i}^{z}}f,\frac{\tilde{P}_{i}^{y}}{\tilde{P}_{i}^{z}}f)\in\mathbb{R}^{2},$
$p(t)$ can be written as a 2D rational Bézier curve with adjusted weights $w_{i}=\tilde{w}_{i}\tilde{P}_{i}^{z}$:
$p(t)$ $=(\frac{\tilde{x}(t)}{\tilde{z}(t)}f,\frac{\tilde{y}(t)}{\tilde{z}(t)}f)$
$=(\frac{\sum_{i=0}^{n}B_{i,n}(t)\tilde{w}_{i}\tilde{P}_{i}^{x}}{\sum_{j=0}^{n}B_{j,n}(t)\tilde{w}_{j}\tilde{P}_{j}^{z}}f,\frac{\sum_{i=0}^{n}B_{i,n}(t)\tilde{w}_{i}\tilde{P}_{i}^{y}}{\sum_{j=0}^{n}B_{j,n}(t)\tilde{w}_{j}\tilde{P}_{j}^{z}}f)$
$=(\sum_{i=0}^{n}\frac{B_{i,n}(t)(\tilde{w}_{i}\tilde{P}_{i}^{z})}{\sum_{j=0}^{n}B_{j,n}(t)(\tilde{w}_{j}\tilde{P}_{j}^{z})}\left(\frac{\tilde{P}_{i}^{x}}{\tilde{P}_{i}^{z}}f\right),\sum_{i=0}^{n}\frac{B_{i,n}(t)(\tilde{w}_{i}\tilde{P}_{i}^{z})}{\sum_{j=0}^{n}B_{j,n}(t)(\tilde{w}_{j}\tilde{P}_{j}^{z})}\left(\frac{\tilde{P}_{i}^{y}}{\tilde{P}_{i}^{z}}f\right))$
$=(\sum_{i=0}^{n}\frac{B_{i,n}(t)w_{i}}{\sum_{j=0}^{n}B_{j,n}(t)w_{j}}P_{i}^{x},\sum_{i=0}^{n}\frac{B_{i,n}(t)w_{i}}{\sum_{j=0}^{n}B_{j,n}(t)w_{j}}P_{i}^{y})$
$=(x(t),y(t))$
$=\sum_{i=0}^{n}\frac{B_{i,n}(t)w_{i}}{\sum_{j=0}^{n}B_{j,n}(t)w_{j}}P_{i}.$
Appendix C Closest Distance between a Point and a Rational Bézier Curve
Given a point $q\in\mathbb{R}^{2}$ and a $n$th-degree rational Bézier curve
$p(t)=\frac{\sum_{i=0}^{n}B_{i,n}(t)w_{i}P_{i}}{\sum_{j=0}^{n}B_{j,n}(t)w_{j}},$
we want to find the closest distance between them and the closest point $t^{\star}$ on the curve $p$. The closest distance is defined as $(p(t^{\star})-q)^{2}$, and $t^{\star}$ is defined as:
$t^{\star}=\arg_{t}\min(p(t)-q)^{2}.$
To solve the $t^{\star}$, we take the derivative of the squared distance with respect to $t$ and set it to zero:
$2(p(t)-q)p^{\prime}(t)=\rho(t)=0,$
which equal to solve the roots of a polynomial with degree $23n-$ 2. For the quadratic rational curve and the cubic rational curve, a 4th order polynomial and a 7th order polynomial need to be solved respectively. We provide the detail separately in section C.1 and section C.2.
To solve all real roots of these polynomials, inspired by DiffVG, we solve the polynomial using bisection and the Newton-Raphson method [45]. The Newton-Raphson solver obtains its initial guess for intervals from isolator polynomials [34]. For the polynomial $\rho(t)$ with two adjacent real roots $t_{1}$ and $t_{2}$, given any two lower order other polynomials $b(t)$ and $c(t)$, define
$a(t)=b(t)\rho^{\prime}(t)+c(t)\rho(t),$
where $\rho^{\prime}(t)$ is the derivative, [34] proofs that $a(t)$ or $b(t)$ must has at least one real root in the closed interval $[t_{0},t_{1}]$. Since $\rho(t_{0})=\rho(t_{1})=0$, we know $a(t_{0})a(t_{1})=b(t_{0})b(t_{1})\rho^{\prime}(t_{0})\rho^{\prime}(t_{1})$. Since $t_{0}$ and $t_{1}$ are roots of $\rho$ and $\rho^{\prime}(t_{0})\rho^{\prime}(t_{1})\leq 0$. Thus either $a(t_{0})a(t_{1})\leq 0$ or $b(t_{0})b(t_{1})\leq 0$. Please see the original paper for discussions on multiple roots.
After solving all the real roots of $a$ and $b$ within the $(0,1)$, the intervals for finding the real roots of $\rho$ are then determined. We can then use bisection and the Newton-Raphson method to find all the roots of $\rho$ within these intervals. The key operation is the selection of the polynomials $b(t)$ and $c(t)$, [34] prefer to find the $b(t)$ and $c(t)$ which follow $\mathrm{Degree}(a)+\mathrm{Degree}(b)=\mathrm{Degree}(\rho)-1$. In the upcoming paragraphs, we will discuss the cases of the 4th order polynomial and the 7th order polynomial in section C.3 and section C.4.
C.1 Closest Distance of Quadratic Rational Curve Case
The quadratic rational curve $p(t)$ can be reorganized using variables $A$ through $F$ as follows:
$p(t)$ $=\frac{(1-t)^{2}w_{0}P_{0}+2(1-t)tw_{1}P_{1}+t^{2}w_{2}P_{2}}{(1-t)^{2}w_{0}+2(1-t)tw_{1}+t^{2}w_{2}}=\frac{At^{2}+Bt+C}{Dt^{2}+Et+F},$
we can further calculate $p(t)-q$ and $p^{\prime}(t)$ as:
$p(t)-q=\frac{(A-Dq)t^{2}+(B-Eq)t+(C-Fq)}{Dt^{2}+Et+F}$
$p^{\prime}(t)=\frac{(AE-BD)t^{2}+(2AF-2CD)t+(BF-CE)}{(Dt^{2}+Et+F)^{2}}$
By reorganizing them using variables $a$ through $f$:
$p(t)-q=\frac{at^{2}+bt+c}{Dt^{2}+Et+F}$
$p^{\prime}(t)=\frac{dt^{2}+et+f}{(Dt^{2}+Et+F)^{2}},$
we can rewrite the target polynomial $\rho(t)$ as below:
$\rho(t)$ $=2(p(t)-q)p^{\prime}(t)=2\frac{(at^{2}+bt+c)(dt^{2}+et+f)}{(Dt^{2}+Et+F)^{3}},$
whose numerator is a 4th order polynomial in the variable $t$. Note that the denominator of $\rho(t)$ is strictly positive, due to the definition of the rational Bézier curve. Therefore, finding the real roots of the equation $\rho(t)=0$ is equivalent to finding the real roots of the 4th order polynomial in the numerator.
C.2 Closest Distance of Cubic Rational Curve Case
The cubic rational curve $p(t)$ can be reorganized using variables $A$ through $H$ as follows:
$p(t)$ $=\frac{(1-t)^{3}w_{0}P_{0}+3(1-t)^{2}tw_{1}P_{1}+3(1-t)t^{2}w_{2}P_{2}+t^{3}w_{3}P_{3}}{(1-t)^{3}w_{0}+3(1-t)^{2}tw_{1}+3(1-t)t^{2}w_{2}+t^{3}w_{3}}$
$=\frac{At^{3}+Bt^{2}+Ct+D}{Et^{3}+Ft^{2}+Gt+H},$
we can further calculate $p(t)-q$ and $p^{\prime}(t)$ as:
$p(t)-q=\frac{(A-Eq)t^{3}+(B-Fq)t^{2}+(C-Gq)t+(D-Hq)}{Et^{3}+Ft^{2}+Gt+H}$
$p^{\prime}(t)=\frac{3A^{2}+2Bt+C}{Et^{3}+Ft^{2}+Gt+H}-\frac{(3Et^{2}+2Ft+G)(At^{3}+Bt^{2}+Ct+D)}{(Et^{3}+Ft^{2}+Gt+H)^{2}}$
By reorganizing them using variables $a$ through $i$:
$p(t)-q=\frac{at^{3}+bt^{2}+ct+d}{Et^{3}+Ft^{2}+Gt+H}$
$p^{\prime}(t)=\frac{et^{4}+ft^{3}+gt^{2}+ht+i}{(Et^{3}+Ft^{2}+Gt+H)^{2}},$
we can write the target polynomial $\rho(t)$ as below:
$\rho(t)$ $=2(p(t)-q)p^{\prime}(t)=2\frac{(at^{3}+bt^{2}+ct+d)(et^{4}+ft^{3}+gt^{2}+ht+i)}{(Et^{3}+Ft^{2}+Gt+H)^{3}},$
whose numerator is a 7th order polynomial in the variable $t$. Note that the denominator of $\rho(t)$ is strictly positive, due to the definition of the rational Bézier curve. Therefore, finding the real roots of the equation $\rho(t)=0$ is equivalent to finding the real roots of the 7th order polynomial in the numerator.
C.3 Isolator Polynomials of the 4th order Polynomial Case
We aim to isolate the real roots of the 4th order polynomial $\rho(t)$ by the roots of two lower order polynomials $a(t)$ and $b(t)$. To keep $\mathrm{Degree}(a)+\mathrm{Degree}(b)=\mathrm{Degree}(\rho)-1$, we select $a(t)$ as the 2nd order polynomial and $b(t)$ as the linear polynomial.
Define $\rho(t)$ and $\rho^{\prime}(t)$ as follows:
$\rho(t)$ $=t^{4}+Bt^{3}+Ct^{2}+Dt+E$
$\rho^{\prime}(t)$ $=4t^{3}+3Bt^{2}+2Ct+D,$
and we know $a(t)=c(t)\rho(t)-b(t)\rho^{\prime}(t).$ Let $c(t)=1$, $a(t)$ can be written as $a(t)=\rho(t)-b(t)\rho^{\prime}(t)$ By doing a long division between $\rho$ and $\rho^{\prime}$, we obtain the isolator polynomials as follows:
$a(t)$ $=(\frac{C}{2}-\frac{3B^{2}}{16})t^{2}+(\frac{3D}{4}-\frac{CB}{8})t+(e-\frac{DB}{16})$
$b(t)$ $=\frac{1}{4}t+\frac{B}{16}$
C.4 Isolator Polynomials of the 7th order Polynomial Case
We aim to isolate the real roots of the 7th order polynomial $\rho(t)$ by the roots of two lower order polynomials $a(t)$ and $b(t)$. To keep $\mathrm{Degree}(a)+\mathrm{Degree}(b)=\mathrm{Degree}(\rho)-1$, we select $a(t)$ and $b(t)$ both be the cubic polynomial.
Define $\rho(t)$ and $\rho^{\prime}(t)$ as follows:
$\rho(t)$ $=t^{7}+p_{6}t^{6}+....+p_{1}t+p_{0}$
$\rho^{\prime}(t)$ $=7t^{6}+6p_{6}t^{5}+....+p_{1},$
and we know $a(t)=c(t)\rho(t)-b(t)\rho^{\prime}(t).$ We wish $a$ and $b$ to both be cubic polynomials. Note that $b(t)p^{\prime}(t)$ is a 9th order polynomial, so we assume $c(t)$ to be a 2nd order polynomial to ensure that $c(t)p(t)$ is a 9th order polynomial too. Let $b(t)=(t^{3}+Ct^{2}+Dt+E)$ and $c(t)=(7t^{2}+At+B)$, we can obtain two 9th order polynomials, and their difference a(x) is a cubic polynomial:
$c(t)\rho(t)$ $=(t^{7}+p_{6}t^{6}+....+p_{1}t+p_{0})(7t^{2}+At+B)$
$b(t)\rho^{\prime}(t)$ $=(7t^{6}+6p_{6}t^{5}+....+p_{1})(t^{3}+Ct^{2}+Dt+E).$
By setting the coefficients of the 8th to 4th order terms of $a(t)$ to 0, we can obtain a system of 5 linear equations in variables A through E. By solving this system of equations, we can determine the values of variables A through E. We can then proceed to find $b(t)$, $c(t)$, and $a(t)$. Please refer to our code for the specific details.
Appendix D More High Quality Results
D.1 More Results of the Text-to-3D Sketch Task
Figure 10: More results of the text-to-3D sketch task.
D.2 More Results of the Image-to-3D Sketch Task
Figure 11: More results of the image-to-3D sketch task
D.3 More 3D Sketch Results Rendered using Blender [9]
Figure 12: More rendered 3D sketch results.
Figure 13: More rendered 3D sketch results.